OpenStudy (nudeoua):

Factoring by grouping? Would someone be able to explain how to do that? With examples? TIA

OpenStudy (abdullahm):

Factor by grouping is when you have 4 terms. You have to find the greatest common factor from the two first terms and from the last two terms. You may have to rearrange the terms so that way when you factor out the greatest common factor from the two pair of terms, so that way you're left with the same expression. Like in the first example, it is (x + y). If you don't get the same expression when you factor something out from it, then you cannot proceed to the next step, and that means the equation is unfactorable. Example 1. ax + bx + ay + by = (ax + bx) + (ay + by) = x(a + b) + y (a + b) = (a+ b)(x + y) Examale 2. (I just rearranged the terms in example 1.) ax + ay + bx + by = (ax + ay) + (bx + by) = a(x + y) + b(x + y) = (a + b)(x + y)

OpenStudy (abdullahm):

So, now, that you know how to factor by grouping, can you help me factor this? x^2 + 5x + 3x + 15 Don't try to add the like terms. Just factor it.

OpenStudy (abdullahm):

If you don't understand what I've said, maybe you can understand something from here: http://www.coolmath.com/algebra/04-factoring/07-factor-by-grouping-01

jhonyy9 (jhonyy9):

@nudeoua so i will try help you - in case of your exercise how can using what have wrote to you above Abdull. ?

jhonyy9 (jhonyy9):

x^2 +5x +3x+15 = by grouping from x^2 +5x what can you factorizing out ?

jhonyy9 (jhonyy9):

what is common factor ?

jhonyy9 (jhonyy9):

the x - so than you factoriz it out what will get ?

jhonyy9 (jhonyy9):

x(x+5) yes ? so the second group please try solve it alone 3x+15 = ?

jhonyy9 (jhonyy9):

moment please - i come back 10 minutes

jhonyy9 (jhonyy9):

so 3x +5*3 = ? what is common factor ?

jhonyy9 (jhonyy9):

so than you dont like collaborate i will go away - sorry

OpenStudy (mathmale):

@nudeous: Given: x^2 + 5x + 3x + 15 Factor the 3 out of the rightmost two terms, please. Show your result. here.